OpenStudy (anonymous):

For school, Joey wants to make spirit ribbons that are each 6 3/4 inches in length. He has a piece of ribbon that is 97 inches in length. What is the greatest number of spirit ribbons Joey can make from the piece of ribbon?

OpenStudy (anonymous):

what are your choes

OpenStudy (mokeira):



OpenStudy (mokeira):

im not so sure though..is GCD inolved?

OpenStudy (mokeira):

@texaschic101

OpenStudy (texaschic101):

if we do the proportion way... 1/(6 3/4) = x/97 cross multiply 6 3/4x = 97 x = 97/ (6 3/4) x = 97 / (27/4) x = 97 * 4/27 x = 388/27 x = 14.37 rounds to 14